# Characterizing Translation-Invariant Bell Inequalities using Tropical Algebra and Graph Polytopes

### v1

@@ II Preliminaries
Throughout the paper, we will use certain abbreviations and specific notation. The range from 00 to $r-1$ is denoted $[r]:=\{0,\dots,r-1\}$. For $N$-partite systems, we denote the parties by $A^{(i)}$, where $i\in[N]$ indexes the party and we label them cyclically, i.e. $i\mod N$. For a single party, the index $i$ is omitted. The bold font, e.g. $\bm{P}$, $\bm{A}$, $\bm{s}$, etc. refers to multipartite quantities (of $N$ parties in general, but not necessarily).
@@ II.1.1 Local, quantum and non-signaling behaviors
A notable property arising in classical and quantum information theory is the non-signaling (NS) property [31, 1]. Let :=$\mathcal{I}:=\{i_{1},i_{2},\dots,i_{|\mathcal{I}|}\}\subset[N]$ be a subset of parties. We define the marginal $\bm{P}^{(\mathcal{I})}(a_{i_{1}},\dots,a_{i_{|\mathcal{I}|}}|x_{0},\dots,x_{N-1})$ by summing over the output $a_{i}$ of the parties that are not in $\mathcal{I}$. A marginal is said to be NS if it is independent of the inputs of the parties that are not in $\mathcal{I}$. In other words, $\bm{P}^{(\mathcal{I})}(a_{i_{1}},\dots,a_{i_{|\mathcal{I}|}}|x_{0},\dots,x_{N-1})=\bm{P}^{(\mathcal{I})}(a_{i_{1}},\dots,a_{i_{|\mathcal{I}|}}|x_{i_{1}},\dots,x_{i_{|\mathcal{I}|}})$. A distribution is said to be NS if for all subsets $\mathcal{I}\subset[N]$ the marginal distributions $\bm{P}^{(\mathcal{I})}$ are NS. For example, distributions associated to LDS or quantum measurements are NS.
@@ II.1.2 LDS in correlator space
The correlator space is a natural coordinate change of the non-signaling space. One-body correlators  [^1] The term correlator is inspired from well-known CHSH experiment. There, the expectation values $\expectationvalue*{A^{(0)}_{x}A^{(1)}_{y}},x,y\in\{0,1\}$ represent the correlations among the parties, and are called correlators. By analogy, any expectation value will be called a correlator. are defined as $\expectationvalue*{A_{x}}=P(0|x)-P(1|x)$ and two-body correlators as $\expectationvalue{A_{x}\cdot B_{y}}=\sum_{a,b=0}^{1}(-1)^{a+b}{P(a,b|x,y)}$. The generalization to more bodies is straightforward. In total there are $(m+1)^{N}-1$ correlators, which we will use as coordinates on the non-signaling space.
@@ II.1.3 Local polytope in correlator space
A coefficient vector $\bm{\alpha}\in\mathbb{R}^{(m+1)^{N}-1}$ defines the linear function $\bm{\alpha}\cdot\bm{q}$, with $\bm{q}\in\mathbb{R}^{(m+1)^{N}-1}$  [^2] In linear algebra terms: $\bm{\alpha}$ lies in the vector space dual to the space $\mathbb{R}^{(m+1)^{N}-1}$ where the local polytope lives.. By virtue of convexity, the minimum value on $\mathcal{L}^{(N,m)}$ is achieved at a vertex:
@@ II.2 Tropical Algebra
Tropical algebra is a mathematical framework for solving discrete optimization problems and analyzing combinatorial structures. Let us begin by defining the arithmetic operations of tropical addition $x\oplus y:=\min(x,y)$ and tropical multiplication $x\odot y:=x+y$. These operations on real numbers plus infinity give rise to the tropical semiring $(\mathbb{R}_{\infty},\oplus,\odot)$, also referred as the min-sum algebra. For a formal exposition, see [25, 44]. Note that the neutral element for tropical addition is $\infty$, and the neutral element for tropical multiplication is 00.
A tropical matrix is a matrix with entries in $\mathbb{R}_{\infty}$; the multiplication of tropical matrices is defined analogously to usual matrix multiplication, with $+$ and $\cdot$ replaced by their tropical analogues:
@@ II.2.2 Kleene plus
Let $F^{\prime}=F-\lambda(F)$ be the normalization of $F$. This substitution ensures that the eigenvalue is $\lambda(F^{\prime})=0$. All cycles in $F^{\prime}$ have non-negative weight, and there exists a cycle of weight 00. Let us define the Kleene plus of the matrix $F^{\prime}$ as follows:
@@
The entry $[F^{\prime+k}]_{i,j}$ gives the minimum weight of all paths of length at most $k$ from node $i$ to $j$ in $\Gamma_{F}$. The sequence $\left(F^{\prime+k}\right)_{k}$ stabilizes at $k=n$ (recall $n$ is the size of the matrix $F$); we will denote the matrix $F^{\prime+n}$ simply by $F^{\prime+}$. Its $(i,j)$-th entry denotes the minimum weight of any path from $i$ to $j$ of any length. In Section VI.2, the Kleene plus is used to classify Bell inequalities.
@@ II.2.3 Computing a tropical eigenvector
One can show that one of the diagonal entries of $F^{\prime+}$ is equal to 00. For such a diagonal entry $[F^{\prime+}]_{j,j}=0$, let $v=F^{\prime+}_{\cdot j}$ be the $j$-th column vector of the matrix $F^{\prime+}$. One can show that $v$ is a tropical eigenvector of $F$ with tropical eigenvalue $\lambda$ [25]. Alternatively, as explained in [46], computing an eigenvector of $F$ is a single source shortest path problem, which can be solved in $O(nm)$ time, where $m$ is again the number of edges in $\Gamma_{F}$. The tropical eigenvector is essential in constructing the critical graph shown in Figure 1.
@@ II.2.4 Critical graph
Figure 1: Weighted directed graph $\Gamma_{F}$ (gray color) and the critical graph $\Gamma_{F}^{\operatorname{crit}}$ (red color). The critical graph $\Gamma_{F}^{\operatorname{crit}}$ is strongly connected and forms a subgraph of $\Gamma_{F}$. All the closed paths on $\Gamma_{F}^{\operatorname{crit}}$ are critical and their normalized weight is $-2$, e.g. the closed path $(1,2,3)$.
@@ II.2.5 Sequence of tropical powers and cyclicity
for the tropical eigenvalue, which is reminiscent of the iterated power algorithm for computing the largest eigenvalue of a matrix (dividing by $N$ is a “tropical $N$-th root”) [48, 27]. In fact, infinitely many 333namely, the entries $\operatorname{tropTr}(F^{\odot N})/N$ for which $F$ contains a minimal weight cycle of length $N$ of the entries in the sequence $(\operatorname{tropTr}(F^{\odot N})/N)_{N=1,2,\dots}$ are already equal to the limit $\lambda(F)$. The other entries can be larger, but they will converge to $\lambda(F)$ when $N\to\infty$. We will see this pattern again when considering Bell inequalities for TI systems with a finite number of sites. One can also consider the sequence of matrices $(F^{\odot N})_{N=1,2,\ldots}$. According to tropical algebra [48, 27], if $F$ is an irreducible matrix, there exist two integers $\sigma$ and $N_{0}$, such that
@@
In the above example, the cyclicity $\sigma(F)$ is $\gcd\{2,3,4\}=1$ since the critical graph $\Gamma_{F}^{\operatorname{crit}}$ is strongly connected and has simple cycles of length $2$, $3$ and $4$. Additionally, $N_{0}$ in Equation 18 is $2$.
@@ III.2 Optimizing a Bell inequality using tropical algebra
In this section, we use tropical algebra to find the classical bound of TINN Bell inequalities. Given coefficients $\bm{\alpha}\in\mathbb{R}^{m+m^{2}}$ for a Bell inequality in the TINN subspace, the associated linear function is given by
@@
The use of tropical algebra reveals a strong connection between Bell inequalities and graph theory. One of the main applications of tropical algebra is finding the shortest path in a weighted directed graph. Let’s consider the graph $\Gamma_{F}$ with nodes labeled by $s\in\mathcal{S}_{\text{LDS}}$ and whose adjacency matrix is  $F$. As explained in Section II.2, the shortest path in $k$ steps from node $s$ to node $t$ in $\Gamma_{F}$ is given by $[F^{\odot k}]_{st}$. Therefore the classical bound of the linear function $I_{\bm{\alpha}}$ can be seen as the shortest normalized closed path of length $N$ in the graph $\Gamma_{F}$.
@@ III.3 The projected local polytope from cycles in the complete graph
Figure 5: Representation of a $4$-node complete directed graph with self-loops. Each node $s\in\mathcal{S}_{\text{LDS}}$ corresponds to an LDS in the $m=2$ scenario. The two-body nearest-neighbor correlator vector $\expectationvalue*{\vec{\bm{A}}{{}_{[R=1]}}}_{(s,t)}$ can be computed using Equation 25. By averaging $\expectationvalue*{\vec{\bm{A}}{{}_{[R=1]}}}_{(s,t)}$ along the closed path in red, we get the projection of the vertex of $\mathcal{L}^{(5,2)}$ associated to the strategy $\bm{s}=(0,0,1,3,1)$.
@@ IV.1 TI-2 projection
[…]
It is then straightforward to express the $N$-partite TI-$2$ polytope:
@@ IV.2 Optimizing TI-$2$ Bell inequalities using tropical algebra
Figure 6: A visualization of the grouping and elimination procedure to construct 2-variable functions when $N=10$ [18]. The gray circles correspond to the eliminated parties. The parties $3j$ and $13j+$ 1 with matching colors are paired to form a new super-party $j$. Terms of the tripartite function $I_{\bm{\alpha}}(s_{i},s_{i+1},s_{i+2})$ containing the party $23j+$ 2 (plain lines of the same color) are summed. The minimization over $s_{3j+2}$ (gray circle) gives the new 2-variable function $\tilde{I}_{\bm{\alpha}}(\bm{s}_{j},\bm{s}_{j+1})$, which can be encoded in a matrix $H$. It turns out that $H=F^{\odot 3}$, where $F$ is the matrix encoding $I_{\bm{\alpha}}(s_{i},s_{i+1},s_{i+2})$ as shown in Equation 45. The tail function $T(\mathbf{s}_{3},\mathbf{s}_{0})$ (purple line) can be encoded as $F^{\odot 1}$. Following Equation 60, the classical bound $\beta=\operatorname{tropTr}(H^{\odot 3}\odot F^{\odot 1})$, where we used $N=3q+r$ with $q=3,r=1$.
@@
$T(\mathbf{s}_{q},\mathbf{s}_{0}):=$ $\sum_{i=3q}^{3q+r}I_{\bm{\alpha}}(s_{i},s_{i+1},s_{i+2}),$ (59)
@@ VI.1 Finding stabilized TINN Bell inequalities
The solutions are $F^{\prime}_{1}$ with $(c_{0},c_{1},c_{2},c_{3},\lambda)=(-1,-1,-1,-1,-1)$ and $F^{\prime}_{2}$ with $(c_{0},c_{1},c_{2},c_{3},\lambda)=(-1,1,1,-1,-1)$. Furthermore, for any real number $\tau$, $\tau F^{\prime}_{1}$ and $\tau F^{\prime}_{2}$ also belong to the set of solutions $\{F^{\prime}:F^{\prime\odot 2}=F^{\prime}\}$. The two inequalities corresponding to $F^{\prime}_{1}$ and $F^{\prime}_{2}$ are $\sum_{i}\expectationvalue*{A^{(i)}_{0}A^{(i+1)}_{0}}\leq N$ and $\sum_{i}\expectationvalue*{A^{(i)}_{1}A^{(i+1)}_{1}}\leq N$. If we allow single-body correlators, we find four additional renormalization-invariant Bell inequalities: $-N\leq\sum_{i}\expectationvalue*{A^{(i)}_{0}}\leq N$ and $-N\leq\sum_{i}\expectationvalue*{A^{(i)}_{1}}\leq N$. All of these inequalities are rather trivial.
@@ VII.2 Dimension analysis of faces with critical graph
Theorem 4 implies that the face defined by $F(\bm{\alpha})$ is spanned by the vectors $\frac{1}{\ell(\bm{c})}\sum_{i=0}^{\ell(\bm{c})-1}\expectationvalue*{\vec{\bm{A}}{{}_{[R=2]}}}_{(c_{i},c_{i+1})}$, where $\bm{c}=(c_{0},c_{1},\dots)$ runs over the simple cycles the critical graph $\Gamma_{F(\bm{\alpha})}^{\operatorname{crit}}$. However, if we only want to compute the dimension of the face, we can do this without having to list all the cycles. For any face $\mathcal{H}_{\bm{\alpha}}$, we can consider the vector space $\operatorname{span}(\mathcal{H}_{\bm{\alpha}})$ spanned by the face. Its dimension is given by $\dim(\operatorname{span}(\mathcal{H}_{\bm{\alpha}}))=\dim(\mathcal{H}_{\bm{\alpha}})+1$ 555 For this to be true we need to ensure that the all zeroes vector doesn’t lie on the affine span of the face, otherwise $\dim(\mathcal{H}_{\bm{\alpha}})=\dim(\operatorname{span}(\mathcal{H}_{\bm{\alpha}}))$. For our polytope $\dim(\mathcal{L}^{(*,m)}_{{\Pi_{2}}})$, one can use the symmetries from Appendix C to argue that the all zeroes vector lies in the interior, and therefore not on the affine span of any face. , in other words the Bell inequality given by $\bm{\alpha}$ is tight if and only if $\dim(\operatorname{span}(\mathcal{H}_{\bm{\alpha}}))=m+Rm^{2}$.

### v2

@@ II Preliminaries
Throughout the paper, we will use certain abbreviations and specific notation. The range from 0 to r1r-1 is denoted $r]:=\{0,\dots,r-1\}$. For $N$-partite systems, we denote the parties by $A^{(i)}$, where $i\in[N]$ indexes the party and we label them cyclically, i.e. $i\mod N$. For a single party, the index $i$ is omitted. The bold font, e.g. $\bm{P}$, $\bm{A}$, $\bm{s}$, etc. refers to multipartite quantities (of $N$ parties in general, but not necessarily).
@@ II.1.1 Local, quantum and non-signaling behaviors
A notable property arising in classical and quantum information theory is the non-signaling (NS) property [31, 1]. Let $\mathcal{I}:=\{i_{1},i_{2},\dots,i_{|\mathcal{I}|}\}\subset[N]$ be a subset of parties. We define the marginal $\bm{P}^{(\mathcal{I})}(a_{i_{1}},\dots,a_{i_{|\mathcal{I}|}}|x_{0},\dots,x_{N-1})$ by summing over the output $a_{i}$ of the parties that are not in $\mathcal{I}$. A marginal is said to be NS if it is independent of the inputs of the parties that are not in $\mathcal{I}$. In other words, $\bm{P}^{(\mathcal{I})}(a_{i_{1}},\dots,a_{i_{|\mathcal{I}|}}|x_{0},\dots,x_{N-1})=\bm{P}^{(\mathcal{I})}(a_{i_{1}},\dots,a_{i_{|\mathcal{I}|}}|x_{i_{1}},\dots,x_{i_{|\mathcal{I}|}})$. A distribution is said to be NS if for all subsets $\mathcal{I}\subset[N]$ the marginal distributions $\bm{P}^{(\mathcal{I})}$ are NS. For example, distributions associated to LDS or quantum measurements are NS.
@@ II.1.2 LDS in correlator space
The correlator space is a natural coordinate change of the non-signaling space. One-body correlators are defined as $\expectationvalue*{A_{x}}=P(0|x)-P(1|x)$ and two-body correlators as $\expectationvalue{A_{x}\cdot B_{y}}=\sum_{a,b=0}^{1}(-1)^{a+b}{P(a,b|x,y)}$. The generalization to more bodies is straightforward. In total there are $(m+1)^{N}-1$ correlators, which we will use as coordinates on the non-signaling space.
@@ II.1.3 Local polytope in correlator space
A coefficient vector $\bm{\alpha}\in\mathbb{R}^{(m+1)^{N}-1}$ defines the linear function $\bm{\alpha}\cdot\bm{q}$, with $\bm{q}\in\mathbb{R}^{(m+1)^{N}-1}$. By virtue of convexity, the minimum value on $\mathcal{L}^{(N,m)}$ is achieved at a vertex:
@@ II.2 Tropical Algebra
Tropical algebra is a mathematical framework for solving discrete optimization problems and analyzing combinatorial structures. Let us begin by defining the arithmetic operations of tropical addition $x\oplus y:=\min(x,y)$ and tropical multiplication $x\odot y:=x+y$. These operations on real numbers plus infinity give rise to the tropical semiring $(\mathbb{R}_{\infty},\oplus,\odot)$, also referred as the min-sum algebra. For a formal exposition, see [25, 42, 43]. Note that the neutral element for tropical addition is $\infty$, and the neutral element for tropical multiplication is 0.
A tropical matrix is a matrix with entries in $\mathbb{R}_{\infty}$; the multiplication of tropical matrices is defined analogously to usual matrix multiplication, with ++ and $\cdot$ replaced by their tropical analogues:
@@ II.2.2 Kleene plus
Let $F^{\prime}=F-\lambda(F)$ be the normalization of $F$. This substitution ensures that the eigenvalue is $\lambda(F^{\prime})=0$. All cycles in $F^{\prime}$ have non-negative weight, and there exists a cycle of weight 0. Let us define the Kleene plus of the matrix $F^{\prime}$ as follows:
@@
The entry $F^{\prime+k}]_{i,j}$ gives the minimum weight of all paths of length at most $k$ from node $i$ to $j$ in $\Gamma_{F}$. The sequence $\left(F^{\prime+k}\right)_{k}$ stabilizes at $k=n$ (recall $n$ is the size of the matrix $F$); we will denote the matrix $F^{\prime+n}$ simply by $F^{\prime+}$. Its $(i,j)$-th entry denotes the minimum weight of any path from $i$ to $j$ of any length. In Section VI.2, the Kleene plus is used to classify Bell inequalities.
@@ II.2.3 Computing a tropical eigenvector
One can show that one of the diagonal entries of $F^{\prime+}$ is equal to 0. For such a diagonal entry $F^{\prime+}]_{j,j}=0$, let $v=F^{\prime+}_{\cdot j}$ be the $j$-th column vector of the matrix $F^{\prime+}$. One can show that $v$ is a tropical eigenvector of $F$ with tropical eigenvalue $\lambda$ [25]. Alternatively, as explained in [45], computing an eigenvector of $F$ is a single source shortest path problem, which can be solved in $O(nm)$ time, where $m$ is again the number of edges in $\Gamma_{F}$. The tropical eigenvector is essential in constructing the critical graph shown in Figure 1.
@@ II.2.4 Critical graph
Figure 1: Weighted directed graph $\Gamma_{F}$ (gray color) and the critical graph $\Gamma_{F}^{\operatorname{crit}}$ (red color). The critical graph $\Gamma_{F}^{\operatorname{crit}}$ is strongly connected and forms a subgraph of $\Gamma_{F}$. All the closed paths on $\Gamma_{F}^{\operatorname{crit}}$ are critical and their normalized weight is 2-2, e.g. the closed path $(1,2,3)$.
@@ II.2.5 Sequence of tropical powers and cyclicity
for the tropical eigenvalue, which is reminiscent of the iterated power algorithm for computing the largest eigenvalue of a matrix (dividing by $N$ is a “tropical $N$-th root”) [47, 27]. In fact, infinitely many of the entries in the sequence $(\operatorname{tropTr}(F^{\odot N})/N)_{N=1,2,\dots}$ are already equal to the limit $\lambda(F)$. More precisely, the entries $\operatorname{tropTr}(F^{\odot N})/N$ for which $F$ contains a minimal weight cycle of length $N$. The other entries can be larger, but they will converge to $\lambda(F)$ when $N\to\infty$. We will see this pattern again when considering Bell inequalities for TI systems with a finite number of sites. One can also consider the sequence of matrices $(F^{\odot N})_{N=1,2,\ldots}$. According to tropical algebra [47, 27], if $F$ is an irreducible matrix, there exist two integers $\sigma$ and $N_{0}$, such that
@@
In the above example, the cyclicity $\sigma(F)$ is $\gcd\{2,3,4\}=1$ since the critical graph $\Gamma_{F}^{\operatorname{crit}}$ is strongly connected and has simple cycles of length $2$, 33 and $4$. Additionally, $N_{0}$ in Equation 18 is $2$.
@@ III.2 Optimizing a Bell inequality using tropical algebra
In this section, we use tropical algebra to find the classical bound of TINN Bell inequalities. As explained in Equation 9, a Bell inequality can be expressed by a vector of coefficients $\bm{\alpha}$. In particular, a Bell inequality in the TINN subspace is defined by the coefficients $\bm{\alpha}\in\mathbb{R}^{m+m^{2}}$. The associated linear functional is then given by
@@
The use of tropical algebra reveals a strong connection between Bell inequalities and graph theory. One of the main applications of tropical algebra is finding the shortest path in a weighted directed graph. Let’s consider the graph $\Gamma_{F}$ with nodes labeled by $s\in\mathcal{S}_{\text{LDS}}$ and whose adjacency matrix is  $F$. As explained in Section II.2, the shortest path in $k$ steps from node $s$ to node $t$ in $\Gamma_{F}$ is given by $F^{\odot k}]_{st}$. Therefore the classical bound of the linear function $I_{\bm{\alpha}}$ can be seen as the shortest normalized closed path of length $N$ in the graph $\Gamma_{F}$.
@@ III.3 The projected local polytope from cycles in the complete graph
Figure 5: Representation of a 44-node complete directed graph with self-loops. Each node $s\in\mathcal{S}_{\text{LDS}}$ corresponds to an LDS in the $m=2$ scenario. The two-body nearest-neighbor correlator vector $\expectationvalue*{\vec{\bm{A}}{{}_{[R=1]}}}_{(s,t)}$ can be computed using Equation 25. By averaging $\expectationvalue*{\vec{\bm{A}}{{}_{[R=1]}}}_{(s,t)}$ along the closed path in red, we get the projection of the vertex of $\mathcal{L}^{(5,2)}$ associated to the strategy $\bm{s}=(0,0,1,3,1)$.
@@ IV.1 TI-2 projection
For example, in the case $m=2$, the strategies $(s_{i},s_{i+1},s_{i+2})=(0,0,3)$ give the correlator vector $\expectationvalue*{\vec{\bm{A}}{{}_{[R=2]}}}_{(0,0,3)}=(1/3,1/3\,|\,0,0,0,0\,|\,-1,-1,-1,-1)^{T}$ using Equations 3 and 41. In Appendix A, we explain how to to similarly define the correlator vector $\expectationvalue*{\vec{\bm{A}}{{}_{[R]}}}_{(s_{i},\ldots,s_{i+R})}$ for $R=3,4,\ldots$; see in particular Equation 84.
It is then straightforward to express the $N$-partite TI-22 polytope:
@@ IV.2 Optimizing TI-$2$ Bell inequalities using tropical algebra
Figure 6: A visualization of the grouping and elimination procedure to construct 2-variable functions when $N=10$ [18]. The gray circles correspond to the eliminated parties. The parties $3j$ and $3j+1$ with matching colors are paired to form a new super-party $j$. Terms of the tripartite function $I_{\bm{\alpha}}(s_{i},s_{i+1},s_{i+2})$ containing the party $3j+2$ (plain lines of the same color) are summed. The minimization over $s_{3j+2}$ (gray circle) gives the new 2-variable function $\tilde{I}_{\bm{\alpha}}(\bm{s}_{j},\bm{s}_{j+1})$, which can be encoded in a matrix $H$. It turns out that $H=F^{\odot 3}$, where $F$ is the matrix encoding $I_{\bm{\alpha}}(s_{i},s_{i+1},s_{i+2})$ as shown in Equation 45. The tail function $T(\mathbf{s}_{3},\mathbf{s}_{0})$ (purple line) can be encoded as $F^{\odot 1}$. Following Equation 60, the classical bound $\beta=\operatorname{tropTr}(H^{\odot 3}\odot F^{\odot 1})$, where we used $N=3q+r$ with $q=3,r=1$.
@@
$T(\mathbf{s}_{q},\mathbf{s}_{0})$= $\sum_{i=3q}^{3q+r}I_{\bm{\alpha}}(s_{i},s_{i+1},s_{i+2}),$ (59)
@@ VI.1 Finding stabilized TINN Bell inequalities
The solutions are $F^{\prime}_{1}$ with $(c_{0},c_{1},c_{2},c_{3},\lambda)=(-1,-1,-1,-1,-1)$ and $F^{\prime}_{2}$ with $(c_{0},c_{1},c_{2},c_{3},\lambda)=(-1,1,1,-1,-1)$. Furthermore, for any real number $\tau$, $\tau F^{\prime}_{1}$ and $\tau F^{\prime}_{2}$ also belong to the set of solutions $\{F^{\prime}:F^{\prime\odot 2}=F^{\prime}\}$. The two inequalities corresponding to $F^{\prime}_{1}$ and $F^{\prime}_{2}$ are $\sum_{i}\expectationvalue*{A^{(i)}_{0}A^{(i+1)}_{0}}\leq N$ and $\sum_{i}\expectationvalue*{A^{(i)}_{1}A^{(i+1)}_{1}}\leq N$. If we allow single-body correlators, we find four additional renormalization-invariant Bell inequalities: $N\leq\sum_{i}\expectationvalue*{A^{(i)}_{0}}\leq N$ and $N\leq\sum_{i}\expectationvalue*{A^{(i)}_{1}}\leq N$. All of these inequalities are rather trivial.
@@ VII.2 Dimension analysis of faces with critical graph
Theorem 4 implies that the face defined by $F(\bm{\alpha})$ is spanned by the vectors $\frac{1}{\ell(\bm{c})}\sum_{i=0}^{\ell(\bm{c})-1}\expectationvalue*{\vec{\bm{A}}{{}_{[R=2]}}}_{(c_{i},c_{i+1})}$, where $\bm{c}=(c_{0},c_{1},\dots)$ runs over the simple cycles the critical graph $\Gamma_{F(\bm{\alpha})}^{\operatorname{crit}}$. However, if we only want to compute the dimension of the face, we can do this without having to list all the cycles. For any face $\mathcal{H}_{\bm{\alpha}}$, we can consider the vector space $\operatorname{span}(\mathcal{H}_{\bm{\alpha}})$ spanned by the face. Its dimension is given by $\dim(\operatorname{span}(\mathcal{H}_{\bm{\alpha}}))=\dim(\mathcal{H}_{\bm{\alpha}})+1$. This is true as long as all zeroes vector do not lie on the affine span of the face (otherwise $\dim(\mathcal{H}_{\bm{\alpha}})=\dim(\operatorname{span}(\mathcal{H}_{\bm{\alpha}}))$. For the polytope $\mathcal{L}^{(*,m)}_{{\Pi_{2}}}$, symmetries (see Appendix C) can be used to argue that the all zeroes vector lies in the interior, and therefore not on the affine span of any face. In other words the Bell inequality given by $\bm{\alpha}$ is tight if and only if $\dim(\operatorname{span}(\mathcal{H}_{\bm{\alpha}}))=m+Rm^{2}$.